Given the following data for heats of reaction

N2(g) + 3H2(g) ----> 2NH3(g)  \DeltaH = -91.8 kJ

C(graphite) + 2H2(g) -------> CH4(g)\DeltaH = -74.9kJ

H2(g) + 2C(graphite) + N2(g) --------> 2HCN(g)  \DeltaH = 270.3 kJ

Calculate \Delta H for the reaction used to make HCN

CH4(g) + NH3(g) --------> HCN(g) + 3H2(g)
Answer #1

Lets number the reaction as 1, 2, 3, 4 from top to bottom
required reaction should be written in terms of other reaction
This is Hess Law
required reaction can be written as:
reaction 4 = -0.5 * (reaction 1) -1 * (reaction 2) +0.5 * (reaction 3)

So, ΔHo rxn for required reaction will be:
ΔHo rxn = -0.5 * ΔHo rxn(reaction 1) -1 * ΔHo rxn(reaction 2) +0.5 * ΔHo rxn(reaction 3)
= -0.5 * (-91.8) -1 * (-74.9) +0.5 * (270.3)
= 256.0 KJ
Answer: 256.0 KJ